OpenStudy (anonymous):

How do I integrate e^6x?

OpenStudy (anonymous):

\[\int\limits e^{6x}\] is this your problem?

OpenStudy (anonymous):

@Dizziedee

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok wait...

OpenStudy (anonymous):

\[\int\limits e^{6x} dx\] let 6x = t Therefore \[6\frac{dx}{dt} = 1 \rightarrow dx = \frac{1}{6}dt\] thus \[\int\limits e^{6x} dx\] becomes \[\int\limits e^{6x} dx =\int\limits\limits e^{t} \frac{1}{6}dt = \frac{1}{6} \int\limits\limits e^{t} dt \] \[= \frac{1}{6} e^{t} = \frac{1}{6} e^{6x} \]( substituting t = 6x ) @Dizziedee